OpenStudy (anonymous):

Use logarithmic differentiation to find dy/dx if y = (x^3 - )^5 / e^7x sin (2x). the given answer = (x^3-1)^5/e^7x sin (2x) [ 15x^2 /x^3-1 -7 - 2 cot 2(2x)]

OpenStudy (anonymous):

you have to use quotient rule

OpenStudy (anonymous):

I deserve a medal!!!!!!!!!!!!!11

OpenStudy (anonymous):

i have try to do it but how to remove e^7x

OpenStudy (kainui):

No, don't use the quotient rule since you were asked to use logarithmic differeniation.

OpenStudy (anonymous):

forget whatever i said

OpenStudy (anonymous):

it okay KissMyAxe

OpenStudy (kainui):

Logarithm differentiation means to take the logarithm of both sides of the equation and then take the derivative. Here's an example: \[\Large y=\frac{(x+7)^6}{ e^{2x}} \\ \Large \ln y = 6 \ln (x+7) -2x \] See, through using log rules you can split it apart, then take the derivative. Then you'll have \[\Large \frac{y'}{y}=6\frac{1}{x+7}-2\] But we want y' so we multiply y to both sides and then plug it in from earlier! Done! It's ok @KissMyAxe =)

OpenStudy (anonymous):

I get it @Kainui

OpenStudy (anonymous):

In y = In (x^3-1)^5 e^(-7x) / In (sin 2x)^1 ln y = 5 ln (x^3-1) - 7x - 1 ln ( sin 2x) (1/y) dy/dx = 5 (3x^2) /(x^3-1) - 7 - (1 /sin 2x) cos (2x) (2) (1/y) dy/dx = 15x^2 /(x^3-1) - 7 - 2 (cos 2x / sin 2x) (1/y) dy/dx = 15x^2 /(x^3-1) - 7 - 2 cot(2x) dy/dx = y [15x^2 /(x^3-1) - 7 - 2 cot(2x) ] dy/dx = (x^3-1)^5 / e^7x sin(2x) ( 15x^2 /(x^3-1) - 7 - 2 cot(2x) ) tq for helping me